. Guppose that the government instinates a ducer surplus as a boses. Show that n total short-run $$.50-pm-6ilmtax on the DVD-copying indu- try. Assuming that the demand for copied fims s that given in punc, how does this tax a ffect the market equilibrium stry. Why dont mputation of the&. How is the burden of this tax allocated r surplus? opying industry is py five DVDs per D. Each firm must and the per-film between consumers and producers? What is the loss of consumer and producer surplus? h. Show that the loss of producer surplas as a result of this tax is borne completely by the fim studios. Explain your results intuitively on of total indu 9,10 The domestic demand for portable rados ia given by Demand: Q 5,000-100P measured in thousands of radios per year. The dome Supply Q 1S0P curve with r as a where price P w measured in dollars and quandry Q is or copied DVDs is te r oopiod DVDs is tk supply cuve for radios is given by for radi 50-50P a. What is the domestic equilibrium in the porta- n long-run equil am price andquan- many DVD firms 6lm royaky rate ble radio market? b. Suppose portablk radios can be imported at a world price of $10 per radio. If trade were unencumbered, what would the new marke equilibrium be? How many portable radios would be produced domestica lly? How many for copied DVDs ble radios would be imported? If domestk portablk radioproducers succeeded in getting a $5 tariff implemented, how would this change the market equilibrium? How much would be colleted in tariff revenues How much consumer suplus would be trans- fered to domestic producers? What would the deadweight loss from the tariff be? 00-S0P n equilibrium price VDs? How many hat is the per-flm libria in the DVD crease in producer d. Graph your results.
Answer #1

a) Domestic equilibrium

where Qd =Qs

5000 - 100P = 150P

5000 = 250P

P* = 20

Q* = 150*20 = 3000

b) at world price of $10,

Qs = 150*10 = 1500 -- domestic supply

Qd = 5000 - 100*10 = 4000 -- domestic demand

so imports will be 4000 - 1500 = 2500 units.

c) Tariff of $5 means now price will be $15

Qd = 5000 - 100*15 = 3500 units

Qs = 150*15 = 2250 units

Imports = 3500 - 2250 = 1250 units

Tariff revenue = 1250*$5 = $6250

Consumer surplus transferred to domestic producers is the area a ( in the diagram)

area a = (15 - 10)* 1500 + (1/2)*(15 -10)*(2250 - 1500) =$1875

Dead weight loss from the tariff is the two yellow shaded triangle

DWL = (1/2)*(15 - 10)*(2250 - 1500) + (1/2)*(15 - 10)*(4000 - 3500) = $3125

d) Graph

60.00 50.00 40.00 30.00 20.00 10.00tarifrevenu Domestic Demand Domestic Supply 15.00 Pw 0.00 200 0 1000 1500 4000 2000 3000 5000 Quantity